(x-1)^2 = -49 take sqrt x-1=7 or x-1=-7 we now add 1 to each side of equations to solve for x x=8 or x=-6
Answer by Alan3354(69443)

(x-1)^2 = -49
Take sqrt:
x-1 = +/- sqrt(-49) = +/- 7i
x = 1 + 7i
x = 1 - 7i